# Future-as-Label: Scalable Supervision from Real-World Outcomes

### v1

@@ Abstract
Many real-world prediction problems lack labels observable at prediction time, creating a temporal gap between prediction and outcome that yields supervision only after events resolve. To address this setting, we extend reinforcement learning with verifiable rewards to temporally resolved real-world prediction, and use it to train language models to make probabilistic forecasts under causally masked information with retrospective evaluation using proper scoring rules. Supervision is derived solely from post-resolution outcomes, preserving delayed-reward semantics. On real-world forecasting benchmarks, Qwen3-32B trained using Foresight Learning improves Brier score by 27% and halves calibration error relative to its pretrained baseline, and outperforms Qwen3-235B on both constructed future-event prediction tasks and the Metaculus benchmark despite a 7× parameter disadvantage.
@@ 4.1.1 Dataset Statistics and Splits
We intentionally do not construct a validation set: models are trained on all available pre-test data using a fixed training procedure, without early stopping or model selection based on held-out examples. Training data consists of predictions made as of July 1, 2024 through January 30, 2025. Both test sets consist exclusively of predictions made on or after February 1, 2025, ensuring strict temporal separation between training and evaluation. The dataset described in this section is released publicly to support reproducibility and future research.

### v2

@@ Abstract
Time creates free supervision: forecasts about real-world events resolve to verifiable outcomes. The passage of time provides labels that require no annotation. To exploit this structure, we extend reinforcement learning with verifiable rewards to real-world prediction over time. We train language models to make probabilistic forecasts from causally masked information, using proper scoring rules as the reward function once events resolve. Learning is driven entirely by realized outcomes, enabling scalable outcome-based supervision in open-world prediction. On real-world forecasting benchmarks, Qwen3-32B trained using Foresight Learning improves Brier score by 27% and halves calibration error relative to its pretrained baseline, and outperforms Qwen3-235B on both constructed future-event prediction tasks and the Metaculus benchmark despite a 7× parameter disadvantage.
@@
Training data: Hugging Face dataset
Model weights: Hugging Face model repo
Data generation: https://lightningrod.ai/
@@ 4.1.1 Dataset Statistics and Splits
We intentionally do not construct a validation set: models are trained on all available pre-test data using a fixed training procedure, without early stopping or model selection based on held-out examples. Training data consists of predictions made as of July 1, 2024 through January 30, 2025. Both test sets consist exclusively of predictions made on or after February 1, 2025, ensuring strict temporal separation between training and evaluation.
@@ 6 Discussion and Conclusion
7 Data and Model Availability
The trained model, datasets, and data generation platform are publicly available to support reproducibility and future research.
Training data: Hugging Face dataset
Model weights: Hugging Face model repo
Data generation: https://lightningrod.ai/